OpenStudy (anonymous):

find domain of (x+4)/(X^2 -4)

jhonyy9 (jhonyy9):

every numbers without -2 and 2

OpenStudy (anonymous):

how do i express that in interval notation i did figure that out haha...thanks

OpenStudy (anonymous):

(x+4)/(x+2)(x-2) \[D:x \neq2,-2\] \[R: y \neq x\]

jhonyy9 (jhonyy9):

- infinite + infinite without -2 and +2

OpenStudy (anonymous):

how would i express that in interval notation though, would it be: (-infinite, -2]U[2, infinite) ?

OpenStudy (anonymous):

almost, (-infinite, -2)U(2, infinite) -2] says that x can be equal to -2, which it cannot.

OpenStudy (anonymous):

The domain in interval notation is \((-\infty,\infty)-\{-2,2\}\), or \((-\infty,-2) ∪ (-2,2)∪ (2,\infty)\).

OpenStudy (anonymous):

ah i see...thank you

OpenStudy (anonymous):

i also have a question posted about a cubed root, finding domain if anyone wishes to help ha..thanks though guys

jhonyy9 (jhonyy9):

put the good answer please for your help